Solve the following systems of equations:


x + 2y = 6 (1)
x = 8-2 y (2)

Answer by Alan3354(69443)

x + 2y = 6 (1)
x = 8-2y (2) --> x + 2y = 8
--------
x + 2y = 6 (1)
x + 2y = 8 (2)
----------
See a problem?
These are inconsistent, no solution.
Their graphs would be 2 parallel lines.